# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v1

@@ Abstract
The first production release of the CUDACPP plugin for the Madgraph5_aMC@NLO generator, which speeds up matrix element (ME) calculations for leading-order (LO) QCD processes using a data parallel approach on vector CPUs and GPUs, was delivered in October 2024. This has been described in previous publications by the team behind that effort. In this paper, I describe my work on some additional developments providing further optimizations of CUDACPP for GPUs, which I consider ready for inclusion in a new release of the software. The new approach mainly consists in splitting the calculation of the ME, which has been so far performed using a single large GPU kernel, into several smaller kernels. I also take this opportunity to describe more in detail some features of the CUDACPP software that are relevant to these new developments and that have not yet been documented.
@@ 1 Introduction
In this context, a recent progress has been the port and optimization of the MG5aMC software for data parallel processing on graphical processing units (GPUs) and on CPUs with SIMD (Single Instruction Multiple Data) vector instructions. After a five-year development programme, which was initially facilitated by the activities of the HEP Software Foundation (HSF) event generator working group bib:csbs , and which regularly reported its progress through conference proceedings bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 , this project delivered its first production release of the MG5aMC code with GPU and SIMD support for QCD leading-order (LO) processes in October 2024. This has been reported at the CHEP2024 conference bib:chep2024 and in a more recent journal submission bib:zw2025 . MG5aMC is a code-generating framework, largely written in Python, which allows users to generate code, by default using Fortran, for any physics process of their choice. The computational bottleneck of the MG5aMC workflow is the calculation of the matrix element (ME) for each phase space point from the momenta of the initial and final state particles in that event. The main outcome of this project is a new code-generating plugin, which was named “CUDACPP” because the generated physics code uses C++ instead of Fortran to execute the calculation of MEs on vector CPUs, and also includes CUDA extensions to run the code on NVidia GPUs (as well as HIP extensions for AMD GPUs). The new code speeds up the ME computation by performing it for many events at the same time using a data parallel approach bib:hsf2020 ; bib:lhcc2020 . To achieve this, work was also needed in the process-agnostic core MG5aMC event processing algorithm for LO processes, MadEvent, which is written in Fortran and with which the process-specific C++ code generated by CUDACPP is interfaced and linked: the main change is the move from a sequential single-event API to a parallel multi-event API, and the refactoring of some functions to make them stateless and re-entrant. This effort has been the result of a collaboration involving many existing and new contributors of the MG5aMC project, as detailed in the references cited above bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 .
@@
In this article, I describe a few new enhancements of the CUDACPP plugin, specifically targeting more efficient processing on GPUs. These are developments for which the initial brainstorming with some of my colleagues from the development team, notably Olivier Mattelaer who prototyped with me the first concrete changes to the code, took place during the CSCS GPU Hackathon that we attended in Lugano in September 2022 bib:hack2022 . I then invested further effort in this area in February and especially in October 2024 bib:ks2024 after the first CUDACPP release, but had to pause this work again due to other constraints. In September 2025, I was finally able to resume this activity, and I believe that some of these developments are now ready to be considered for their inclusion in a new production release, which is why I found it useful to document them in some detail. The main idea behind all of these developments can be referred to as “kernel splitting”. In short, this consists in replacing the large monolithic kernel currently used by CUDACPP for the calculation of MEs from particle momenta, named sigmaKin, by many smaller kernels, executed either in parallel or sequentially, to achieve a more efficient and scalable use of the GPU. The aim is not only that of possibly achieving a higher peak throughput (in terms of MEs computed per second) with a large GPU grid, i.e. with a large number of events processed in parallel on the GPU during one offloading cycle, but also that of achieving a faster increase of throughput already with smaller GPU grids. An important practical difference between the traditional MG5aMC workflow on CPUs and that using GPUs, in fact, is that the former was designed to handle iterations involving hundreds of events at a time, while the latter typically needs many thousands of events to be efficient: in this context, achieving higher throughputs with smaller grids would provide one way to use more manageable, but still reasonably efficient, event generation jobs.
@@
More in detail, in this paper I describe my work on four sets of kernel splitting developments: (1) helicity streams; (2) color sum splitting into separate GPU kernels; (3) color sum host refactoring with optional BLAS offloading; and (4) Feynman diagram splitting into separate GPU kernels. The performance of each set of developments was tested for different physics processes, software configurations and hardware architectures. While these developments are almost exclusively aimed at improving throughputs on GPUs, the large code refactoring that they imply was also propagated to the vectorised C++ code for SIMD CPUs, as code generated by the CUDACPP plugin is strictly the same for CPUs and and GPUs, and the distinction between the two cases is only done at build time through #ifdef directives. I therefore analysed not only CUDA and HIP performance on NVidia and AMD GPUs, but also C++ performance on vector CPUs using different SIMD configurations. Considering the results of all my tests, my recommendation is to include the first three enhancements in a new production release of CUDACPP, while keeping BLAS offloading as an option that is disabled by default. The fourth development, conversely, results in a significant performance degradation on GPUs, and to a lesser extent also on CPUs, and my recommendation is to keep this as a proof-of-concept implementation in a separate development branch, in view of possible further refinements of this approach.
@@
As these new developments imply a significant refactoring of the existing CUDACPP code base, and rely on internal features that have not yet been documented, I also found it useful to start this article with a brief review of some architectural aspects of the software. In this context, I should mention that the internal design of the CUDACPP plugin, which was largely my own work, is far from perfect and the code could certainly benefit from an extensive cleanup. In particular, over time I have added various switches to support alternative parallel implementations of some software components: I believe that this has been very useful, and in some cases essential, to try out different approaches and eventually converge on well-defined solutions to address specific problems, but I am aware that this has resulted in code that might be more difficult to read, and where some of this flexibility may now represent an unnecessary complication. I hope that this article may help clarify the rationale behind some design choices and provide guidance for possible future developments and/or cleanups. I stress, in any case, that the opinions that I express in this paper are only my own, and may differ from those of some of my colleagues.
@@
The outline of this paper is the following. In Sec. 2, I briefly review some aspects of the CUDACPP software architecture design, focusing on those that are most relevant to kernel splitting. In Sec. 3, I describe in detail the software design and implementation of the various kernel splitting enhancements, and I provide the results and an analysis of various performance tests. In Sec. 4, I give my conclusions and an an outlook for this work.
@@ 2 Software architecture review
The overall software architecture design of the CUDACPP plugin for GPUs and vector CPUs, its evolution over time, its interplay with the MG5aMC framework and many specific implementation choices have already been described in the previous papers bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 from our team and I will not repeat this here. The main point of this section, instead, is to give further information about some software aspects which have not yet been documented and which are relevant to kernel splitting, such as memory layouts and memory access. First, in any case, I will give a very high level overview of the various areas and phases of development on CUDACPP, to put the kernel splitting work into context.
@@ 2.1 Three software areas and development phases
Over time, the focus of developments has gradually shifted from the first to the last of these software areas, even if some work on ME engines and on the integration of C++ and Fortran continued up until the release in October 2024. This is reflected in the documentation provided in our previous papers. What I want to stress here, to put it into context, is that the new kernel splitting work that I describe in Sec. 3 was all done at the level of the ME engines alone, even if its motivation (as is the case for all of the work in the Madgraph on GPU project) lies in the streamlining of the full MG5aMC workflow. I mention this because I believe that, amongst the many previous documents from this project, those describing the software features most relevant to kernel splitting in the ME engines are actually the very first ones, namely the proceedings bib:vchep2021 and presentation bib:vchep2021video from the vCHEP2021 conference.
More specifically, in all of my work on kernel splitting presented in Sec. 3, I only developed code using the standalone application check.exe, where all three main components of an event generation application (random numbers, phase space sampling and ME calculation) are implemented in C++ and CUDA using a data parallel approach. This is because check.exe makes it possible to focus on the optimization of the ME engine without being constrained by the MadEvent Fortran infrastructure, where the legacy phase space sampling and other sequential non-ME components slow down the whole workflow. I routinely built check.exe using the CUDACPP makefile cudacpp.mk, without even building MadEvent or any other Fortran code (which proceeds through a separate makefile that internally delegates the build of the CUDACPP library to cudacpp.mk).
@@
Table 1: Pseudo-code of the ME calculation in the latest CUDACPP, before any kernel splitting changes. Function main represents the check.exe standalone application. Note that sigmaKin is the only GPU kernel (__global__ function) in this workflow. Only a simplified GPU version of the code is shown, using scalar fptype floating point and cxtype complex number types: the actual code that supports both GPUs and SIMD CPUs uses fptype_sv and cxtype_sv data types inside calculate_wavefunction, making it possible to use the same formal code for scalar data and for SIMD vectors (see Ref. bib:vchep2021 for details). The CPU branch of the code for main, encapsulated by an #ifdef and not shown here, uses the Host versions of various classes instead of the Device versions. The pseudo-code for the color sum schematically indicates that this proceeds via a quadratic form using a color matrix, and that color amplitudes are converted from fptype/cxtype floating point precision to a potentially different fptype2/cxtype2 precision (this enables the “mixed” precision mode, where the former is based on double and the latter on float).
@@ 2.2 A closer look at the internals of the ME engine
Consider a physics process where npar is the total number of initial and final state particles. Assume that the calculation of Feynman diagram amplitudes is performed using floating types fptype, which is a typedef for either double or float. Very schematically, as shown in Fig. 1, the ME computational engine of CUDACPP is simply a software component that takes as input an array of particle momenta for many events (i.e. 4*npar fptype values for each of nevt events), and returns as its output an array of matrix elements (i.e. one fptype value for each of nevt events). The calculation in CUDACPP is actually more complex than this, as the ME engine also receives other inputs, such as: an event-by-event running coupling $\alpha_{s}$ (at a scale which is computed from particle momenta using an external Fortran module); two event-by-event random numbers for the selection of one specific helicity and leading color in the generated event (which are also returned as outputs); and an event-by-event channel identifier (for MadEvent single-diagram enhancement, also resulting in additional outputs). However, I will neglect this in the following, except otherwise stated, and focus only on the calculation of output MEs from input momenta.
@@
In the current production version of the CUDACPP plugin, v1.00.02 in MG5aMC v3.6.3, the ME computational engine on GPUs is essentially a single and very large monolithic GPU kernel, named sigmaKin. This performs all of the relevant operations that are needed to compute the output ME from the input momenta for each event (one notable exception being the calculation of other couplings that depend on $\alpha_{s}$, which is delegated to a separate kernel). Very schematically, this is represented in the pseudo-code in Table 1. In particular, sigmaKin performs the following operations for each event: it keeps a running sum of the ME over helicities; it loops on all possible helicities of the external particles (or, more precisely, only on a pre-determined set of “good” helicities whose contribution is non-zero); it adds to the running sum of the ME the contribution from each helicity, by calling a device function calculate_wavefunctions which internally involves for each QCD color flow the calculation of particle and propagator wavefunctions and that of dual amplitudes, followed by the sum over all color flows using a quadratic form based on a color matrix; after the end of the helicity loop, sigmaKin randomly draws a helicity and a color for the generated event based on the ME contributions from individual helicities and colors. The calculation of helicity amplitudes for Feynman diagrams, in particular, is described in detail in Ref. bib:vchep2021 (see Fig. 1 therein); complementary details are also given in Ref. bib:zw2025 . As shown schematically in Table 1, this involves three types of elements to compute: the wavefunctions of external (initial/final state) particles, via functions like VXXXXX; the wavefunctions of internal propagators, e.g. via VVV1P0_1; the dual amplitudes for a given colour flow, e.g. via FFV1_0. The names and roles of these functions in CUDACPP are the same as in the original HELAS bib:helas and ALOHA bib:aloha implementations.
@@
    Memory layout and data access. The specific layout chosen for storing data inside the raw memory buffers allocated in the previous step is encapsulated in a separate set of memory access classes. All these classes have methods named kernelAccess, or some variation of this. On the GPU, these methods take as input the fptype* pointer associated to a raw memory buffer for nevt events, and they return as output a data item for the single event ievt processed by the GPU thread where the code is executed, typically indexed by the identifier of the GPU thread, blockDim.x*blockIdx.x+threadIdx.x. In some cases, these methods have additional parameters: method kernelAccessIp4IparConst in the DeviceAccessMomenta class, for instance, takes two additional parameters ip4 and ipar in order to retrieve a given 4-vector component for a given particle. It is only the memory access classes that, internally, are able to decode the memory layouts of a raw buffer: this is done by interpreting the buffers as one-dimensional C-style array or casting them as multi-dimensional C-style arrays. For most data buffers, a Structure-Of-Array (SOA) or an Array-Of-Structure-Of-Array (AOSOA) layout is chosen, where the data items of a given type for different events are contiguous: as explained in detail in the vCHEP2021 proceedings bib:vchep2021 , this is absolutely essential for SIMD processing on vector CPUs, and it is also beneficial – but not at all a strict requirement – on GPUs to improve performance through coalesced data access. In some cases, such as couplings, the floating point raw arrays include the real and imaginary components of complex data, and it is the responsibility of the data access classes to return outputs with the API of a complex number data type. The case of CPU code, which I will not discuss here in detail as it is less relevant to kernel splitting, includes the additional complication of returning outputs whose data types fptype_sv and cxtype_sv can be scalar (for no-SIMD C++) or SIMD vectors (through compiler vector extensions bib:vchep2021 , hereafter CVEs) of fptype or cxtype data. Concerning the use of memory access classes in other CUDACPP software components, notably the computational methods for helicity amplitudes like VXXXXX, VVV1P0_1 or FFV1_0, I initially chose to implement this as a template parameter: in Table 1, for instance, DeviceAccessMomenta is a template parameter M_ACCESS of VXXXXX. The idea behind the choice of using templates was to allow the flexibility of easily switching between different memory layouts at build time, to compare their performances; in retrospective, the same flexibility may have been achieved with the same memory access classes but without templates. One could now consider removing those templates, although it is not clear to me if this would make the code easier to read and maintain, or faster to build.

### v2

@@ Abstract
The first production release of the CUDACPP plugin for the Madgraph5_aMC@NLO generator, which speeds up matrix element (ME) calculations for leading-order (LO) processes using a data parallel approach on vector CPUs and GPUs, was delivered in October 2024. This has been described in previous publications by the team behind that effort. In this paper, I describe my work on some additional developments and optimizations of CUDACPP, mainly but not exclusively for GPUs. The new approach, which represents a major restructuring of the CUDACPP computational engine, mainly consists in splitting the calculation of the ME, which was previously performed using a single large GPU kernel, into many smaller kernels. A first batch of changes, involving the move to separate “helicity streams” and the optional offloading of QCD color sums to BLAS, also via GPU tensor cores, was recently merged into a new production release of CUDACPP, in collaboration with my colleagues. Since then, I have completed a second batch of changes, which I now also consider ready to be merged in production, involving the possibility to split the calculation into groups of Feynman diagrams in separate source code files. This new feature makes it possible to compute QCD matrix elements for physics processes with a larger number of final state gluons: in particular, I present the first performance results from CUDACPP for the $2\!\rightarrow\!6$ process $gg\!\rightarrow\!t\bar{t}gggg$ on CPUs and GPUs and the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$ on CPUs, which involve over 15k and 230k Feynman diagrams, respectively. I also take this opportunity to describe in detail some features of the CUDACPP software which had not yet been documented, both in the GPU and vector CPU implementations.
@@ 1 Introduction
In this context, a recent progress has been the port and optimization of the MG5aMC software for data parallel processing on graphical processing units (GPUs) and on CPUs with SIMD (Single Instruction Multiple Data) vector instructions. After a five-year development programme, which was initially facilitated by the activities of the HEP Software Foundation (HSF) event generator working group bib:csbs , and which regularly reported its progress through conference proceedings bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 , this project delivered its first production release of the MG5aMC code with GPU and SIMD support for QCD leading-order (LO) processes in October 2024. This has been reported at the CHEP2024 conference bib:chep2024 and in a more recent journal submission bib:zw2025 . MG5aMC is a code-generating framework, largely written in Python, which allows users to generate code, by default using Fortran, for any physics process of their choice. For complex processes, the computational bottleneck of the MG5aMC workflow is the calculation of the matrix element (ME) for each phase space point from the momenta of the initial and final state particles in that event. The main outcome of this project has been a new code-generating plugin, which was named “CUDACPP” because the generated physics code uses C++ instead of Fortran to execute the calculation of MEs on vector CPUs, and also includes CUDA extensions to run the code on NVidia GPUs (as well as HIP extensions for AMD GPUs). The new code speeds up the ME computation by performing it for many events at the same time using a data parallel approach bib:hsf2020 ; bib:lhcc2020 . To achieve this, work was also needed in the process-agnostic core MG5aMC event processing algorithm for LO processes, MadEvent, which is written in Fortran and with which the process-specific C++ code generated by CUDACPP is interfaced and linked: the main change is the move from a sequential single-event API to a parallel multi-event API, and the refactoring of some functions to make them stateless and re-entrant. This effort has been the result of a collaboration involving many existing and new contributors of the MG5aMC project, as detailed in the references cited above bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 .
In this paper, I describe my work on many additional enhancements and optimizations of CUDACPP, mainly but not exclusively targeting GPUs. These are developments for which the initial brainstorming with some of my colleagues from the development team, notably Olivier Mattelaer who prototyped with me the first concrete changes to the code, took place during the CSCS GPU Hackathon that we attended in Lugano in September 2022 bib:hack2022 . I then invested further effort in this area in February and especially in October 2024 bib:ks2024 after the first CUDACPP release, but had to pause this work again due to other constraints. In September 2025, I was finally able to resume this R&D activity, which I now consider essentially complete and which I am documenting in detail in this paper in view of its possible integration in new production versions of the CUDACPP software. A first batch of changes, which I had described in the preprint of a previous version of this paper bib:preprintV1 , has already been merged into a new production release of CUDACPP bib:cudacpp364 , in collaboration with my colleagues. Since then, I have completed a second batch of changes, which were still work in progress at the time of Ref. bib:preprintV1 , but which I now also consider ready to be merged in production.
1.1 Kernel splitting
@@
The main idea behind all of these developments can be referred to as “kernel splitting”. In short, this consists in replacing the large monolithic kernel previously used by CUDACPP for the calculation of MEs from particle momenta, named sigmaKin, by many smaller kernels, executed either in parallel or sequentially, to achieve a more efficient and scalable use of the GPU. The aim is not only that of possibly achieving a higher peak throughput (in terms of MEs computed per second) with a large GPU grid, i.e. with a large number of events processed in parallel on the GPU during one offloading cycle, but also that of achieving a faster increase of throughput already with smaller GPU grids. This was the main focus of my first batch of changes, which have since been released. An important practical difference between the traditional MG5aMC workflow on CPUs and that using GPUs, in fact, is that the former was designed to handle iterations involving hundreds of events at a time, while the latter typically needs many thousands of events to be efficient: in this context, achieving higher throughputs with smaller grids would provide one way to use more manageable, but still reasonably efficient, event generation jobs on GPUs. An additional benefit of kernel splitting, which eventually became the main focus of my second batch of changes, is that smaller kernels are much more manageable by humans and compilers alike (both for CPUs and GPUs), and make it possible to achieve complex calculations that are instead very difficult if not impossible to perform with large monolithic kernels, such as the calculations of QCD matrix elements for physics processes with a large number of final state gluons. In this paper, in particular, I present the first performance results from CUDACPP for the $2\!\rightarrow\!6$ process $gg\!\rightarrow\!t\bar{t}gggg$ on CPUs and GPUs and for the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$ on CPUs, which involve over 15k and over 230k Feynman diagrams, respectively.
ihel0
cudacpp v1.00.02/3.6.3
one single kernel for all helicities (monolithic sigmaKin kernel)
ihel1
one kernel per helicity; one CUDA Stream per helicity
ihel2
two kernels per helicity: color sum & diagrams
QCD
color sum
Feynman
diagrams
ihel3
color sum: one kernel per helicity OR one BLAS per helicity
ihel4
one kernel per diagram in each helicity stream
ihel3p1
cudacpp v1.01.01/3.6.4
color sum: one kernel per helicity OR one single BLAS for all helicities
ihel4p1
one kernel per diagram in each helicity stream; CUDA Graph orchestration
ihel5
one kernel per diagram group in each helicity stream; separate source code files
ihel6p1
(DCDIAG=0) one kernel per diagram group in each helicity stream, exchange wavefunctions in GPU global memory, exchange color amplitudes in GPU global memory OR (DCDIAG=1) one single kernel for all diagram groups in each helicity stream, exchange wavefunctions in GPU registers, exchange color amplitudes in GPU registers
ihel6p2
remove templates from helicity amplitude methods
Figure 1: Schematic representation of the architectural evolution of matrix element calculations in the MG5aMC CUDACPP plugin, in the context of the kernel splitting developments presented in this paper. The steps above the blue dotted line, up to ihel3 and ihel4 included, have been presented in a previous preprint bib:preprintV1 of this paper, while those below it are described here for the first time. The steps in the green boxes, up to ihel3p1 included, have been merged upstream and included in production releases of CUDACPP; the steps in the red boxes, conversely, from ihel4 to ihel6p2 included, are included in a pull request that I recommend for merging upstream. More details about each step are provided in the text and in other figures.
@@
Schematically, in this paper I describe my work on six sets of kernel splitting developments: (1) helicity streams; (2) color sum splitting into separate GPU kernels; (3) color sum host refactoring with optional BLAS offloading; (4) Feynman diagram splitting, with one separate GPU kernel for each diagram; (5) Feynman diagram splitting, with one GPU kernel for each group of diagrams, and separate source code files; (6) Feynman diagram splitting, with a single GPU kernel processing all diagram groups as GPU device functions. A brief description of these steps and of the connection between them is shown in Fig. 1. The performance of each set of developments was tested for different physics processes, software configurations and hardware architectures. While these developments were initially aimed at improving throughputs on GPUs, the large code refactoring that they imply was also propagated to the vectorised C++ code for SIMD CPUs, as code generated by the CUDACPP plugin is strictly the same for CPUs and GPUs, and the distinction between the two cases is only done at build time through #ifdef directives. I therefore analysed not only CUDA and HIP performance on NVidia and AMD GPUs, but also C++ performance on vector CPUs using different SIMD configurations. The possibility to split the ME calculation into groups of Feynman diagrams in different files, in particular, was the main enabling factor for being able to compute complex $2\!\rightarrow\!7$ physics processes on CPUs.
Considering the results of all my tests, my recommendation is to include the full set of these enhancements in new production releases of CUDACPP. This differs from the situation that I had described in my previous preprint bib:preprintV1 , where the diagram splitting work had not yet been completed. A first batch of changes (green boxes in Fig. 1) has already been merged into a new production release, in collaboration with my colleagues: this includes the first three steps of this work, notably helicity streams, the separation of color sums and Feynman diagrams, and the optional BLAS offloading of color sums (which was significantly improved with respect to that described in Ref. bib:preprintV1 ). The second batch of developments (red boxes in Fig. 1), notably including the splitting of Feynman diagrams into different groups and source code files, is in my opinion now also ready to be merged upstream. I note, in particular, that the performance of the code in its default configuration would not be negatively impacted by these changes: many of the features that I have added, in fact, are optional and are only enabled with specific flags at generation time, build time or run time (this includes the BLAS offloading of color sums, the ME splitting into groups of Feynman diagrams, their orchestration via CUDA graphs and the choice to treat diagram groups as separate kernels or as device functions within a single kernel).
1.2 Outline of this paper
@@
As these new developments imply a significant refactoring of the existing CUDACPP codebase, and rely on internal features that have not yet been documented, I also found it useful to start this article with a brief review of some architectural aspects of the software, focusing on those most relevant to the kernel splitting developments, but also covering some details of the SIMD implementation. In this context, I should mention that the internal design of the CUDACPP plugin, which was largely my own work, is far from perfect and the code could certainly benefit from an extensive cleanup. In particular, over time I have added various switches to support alternative implementations of some software components, as I have done again in many of the new developments described in this paper: I believe that this has been very useful, and in some cases essential, to try out different approaches and eventually converge on well-defined solutions to address specific problems, but I am also aware that this results in code that might be more difficult to read, and where some of this flexibility may be regarded as an unnecessary complication. I hope that this article may help clarify the rationale behind some of these design choices and provide some guidance for possible future developments and/or cleanups. I stress, in any case, that the opinions that I express in this paper are only my own, and may differ from those of some of my colleagues.
@@
The outline of this paper is the following. In Sec. 2, I briefly review some aspects of the CUDACPP software architecture design, focusing on those relevant to GPU kernel splitting, but also covering some details of the SIMD implementation for vector CPUs. In Sec. 3, I describe in detail the software design and implementation of my first batch of kernel splitting enhancements, and I provide the results and an analysis of various performance tests: this includes and partially extends the material that I had covered in my previous preprint bib:preprintV1 , with the exception of Feynman diagram splitting. In Sec. 4, I describe the various options I introduced to split the calculation of Feynman diagrams into different groups of diagrams and/or different kernels, and different source code files; I also give the first CUDACPP performance results for physics processes with large numbers of final state gluons, which could be computed thanks to these new features. In Sec. 5, I finally give my conclusions and an an outlook for this work.
@@ 2 Software architecture review
The overall software architecture design of the CUDACPP plugin for GPUs and vector CPUs, its evolution over time, its interplay with the MG5aMC framework and many specific implementation choices have already been described in the previous papers bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 from our team and I will not repeat this here. The main point of this section, instead, is to give further information about some software aspects which have not yet been documented, mainly those relevant to GPU kernel splitting, such as memory layouts and memory access, but also some related to the SIMD implementation for vector CPUs. First, in any case, I will give a very high level overview of the various areas and phases of development on CUDACPP, to put into context the new kernel splitting developments.
@@ 2.1 Three software areas and development phases
Over time, the focus of developments has gradually shifted from the first to the last of these software areas, even if some work on ME engines and on the integration of C++ and Fortran continued up until the release in October 2024. This is reflected in the documentation provided in our previous papers. What I want to stress here, to put it into context, is that the new kernel splitting work that I describe in this paper was all done at the level of the ME engines alone, even if its motivation (as is the case for all of the work in the Madgraph on GPU project) lies in the streamlining of the full MG5aMC workflow. I mention this also because I believe that, amongst the many previous documents from this project, those describing the software features most relevant to kernel splitting in the ME engines are actually the very first ones, namely the proceedings bib:vchep2021 and presentation bib:vchep2021video from the vCHEP2021 conference.
More specifically, in all of the work that I present in this paper, I only developed code using the standalone application check.exe, where all three main components of an event generation application (random numbers, phase space sampling and ME calculation) are implemented in C++ and CUDA using a data parallel approach. This is because check.exe makes it possible to focus on the optimization of the ME engine without being constrained by the MadEvent Fortran infrastructure, where the legacy phase space sampling and other sequential non-ME components slow down the whole workflow. I routinely built check.exe using the CUDACPP makefile cudacpp.mk, without even building MadEvent or any other Fortran code (which proceeds through a separate makefile that internally delegates the build of the CUDACPP library to cudacpp.mk).
@@
Table 1: Pseudo-code of the ME calculation in CUDACPP before any kernel splitting changes (“ihel0”). Function main represents the check.exe standalone application. Note that sigmaKin is the only GPU kernel (__global__ function) in this workflow. Only a simplified GPU version of the code is shown, using scalar fptype floating point and cxtype complex number types: the actual code that supports both GPUs and SIMD CPUs uses fptype_sv and cxtype_sv data types inside calculate_wavefunction, making it possible to use the same formal code for scalar data and for SIMD vectors (see Ref. bib:vchep2021 for details). The CPU branch of the code for main, encapsulated by an #ifdef and not shown here, uses the Host versions of various classes instead of the Device versions. The pseudo-code for the color sum schematically indicates that this proceeds via a quadratic form using a color matrix, and that color amplitudes are converted from fptype/cxtype floating point precision to a potentially different fptype2/cxtype2 precision (this enables the “mixed” precision mode, where the former is based on double and the latter on float).
@@ 2.2 A closer look at the internals of the ME engine
Consider a physics process where npar is the total number of initial and final state particles. Assume that the calculation of Feynman diagram amplitudes is performed using floating point fptype, which is a typedef for either double or float. Very schematically, as shown in Fig. 2, the ME computational engine of CUDACPP is simply a software component that takes as input an array of particle momenta for many events (i.e. 4*npar fptype values for each of nevt events), and returns as its output an array of matrix elements (i.e. one fptype value for each of nevt events). The calculation in CUDACPP is actually more complex than this, as the ME engine also receives other inputs, such as: an event-by-event running coupling $\alpha_{s}$ (at a scale which is computed from particle momenta using an external Fortran module); two event-by-event random numbers for the selection of one specific helicity and leading color in the generated event (which are also returned as outputs); and an event-by-event channel identifier (for MadEvent single-diagram enhancement, also resulting in additional outputs). However, I will neglect this in the following, except otherwise stated, and focus only on the calculation of output MEs from input momenta.
@@
In the production version of the CUDACPP plugin before any kernel splitting changes (hereafter “ihel0”), v1.00.02 in MG5aMC v3.6.3, the ME computational engine on GPUs is essentially a single and very large monolithic GPU kernel, named sigmaKin. This performs all of the relevant operations that are needed to compute the output ME from the input momenta for each event (one notable exception being the calculation of other couplings that depend on $\alpha_{s}$, which is delegated to a separate kernel). Very schematically, this is represented in the pseudo-code in Table 1. In particular, sigmaKin performs the following operations for each event: it keeps a running sum of the ME over helicities; it loops on all possible helicities of the external particles (or, more precisely, only on a pre-determined set of “good” helicities whose contribution is non-zero); it adds to the running sum of the ME the contribution from each helicity, by calling a device function calculate_wavefunctions which internally involves for each QCD color flow the calculation of particle and propagator wavefunctions and that of dual amplitudes, followed by the sum over all color flows using a quadratic form based on a color matrix; after the end of the helicity loop, sigmaKin randomly draws a helicity and a color for the generated event based on the ME contributions from individual helicities and colors. The calculation of helicity amplitudes for Feynman diagrams, in particular, is described in detail in Ref. bib:vchep2021 (see Fig. 1 therein); complementary details are also given in Ref. bib:zw2025 . As shown schematically in Table 1, this involves three types of elements to compute: the wavefunctions of external (initial/final state) particles, via functions like VXXXXX; the wavefunctions of internal propagators, e.g. via VVV1P0_1; the dual amplitudes for a given colour flow (“jamps”), e.g. via FFV1_0. The names and roles of these functions in CUDACPP are the same as in the original HELAS bib:helas and ALOHA bib:aloha implementations.
@@
    Memory layout and data access. The specific layout chosen for storing data inside the raw memory buffers allocated in the previous step is encapsulated in a separate set of memory access classes. All these classes have methods named kernelAccess, or some variation of this. On the GPU, these methods take as input the fptype* pointer associated to a raw memory buffer for nevt events, and they return as output a data item for the single event ievt processed by the GPU thread where the code is executed, typically indexed by the identifier of the GPU thread, blockDim.x*blockIdx.x+threadIdx.x. In some cases, these methods have additional parameters: method kernelAccessIp4IparConst in the DeviceAccessMomenta class, for instance, takes two additional parameters ip4 and ipar in order to retrieve a given 4-vector component for a given particle. It is only the memory access classes that, internally, are able to decode the memory layouts of a raw buffer: this is done by interpreting the buffers as one-dimensional C-style array or casting them as multi-dimensional C-style arrays. For most data buffers, a Structure-Of-Array (SOA) or an Array-Of-Structure-Of-Array (AOSOA) layout is chosen, where the data items of a given type for different events are contiguous: as explained in detail in the vCHEP2021 proceedings bib:vchep2021 , this is absolutely essential for SIMD processing on vector CPUs, and it is also beneficial – but not at all a strict requirement – on GPUs to improve performance through coalesced data access. In some cases, such as couplings, the floating point raw arrays include the real and imaginary components of complex data, and it is the responsibility of the data access classes to return outputs with the API of a complex number data type. The case of CPU code includes the additional complication of returning outputs whose data types fptype_sv and cxtype_sv can be scalar (for no-SIMD C++) or SIMD vectors fptype_v and cxtype_v (through compiler vector extensions bib:vchep2021 , hereafter CVEs) of fptype or cxtype data. Concerning the use of memory access classes in other CUDACPP software components, most notably in the computational methods for helicity amplitudes like VXXXXX, VVV1P0_1 or FFV1_0, I initially chose to implement this as a template parameter: in Table 1, for instance, DeviceAccessMomenta is a template parameter M_ACCESS of VXXXXX. The idea behind the choice of using templates was to allow the flexibility of easily switching between different memory layouts at build time, to compare their performances; in retrospective, the same flexibility may have been achieved with the same memory access classes but without templates [^1] As noted in Sec. 4, eventually I did decide to remove those templates, even if it was not clear to me a priori if this would make the code easier to read and maintain, or faster to build: my main motivation was an attempt to allow some CUDA builds in some configurations of particularly complex $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes, which were otherwise crashing or taking too long. Unfortunately, my tests about the effectiveness of template removal for simplifying CUDA builds were at best inconclusive, as those builds continued to fail; however, removing templates did not have any negative impact on performance elsewhere, and I suggest that this may represent a useful cleanup to merge upstream and integrate in a future production release. .